# CtRL-Sim: Reactive and Controllable Driving Agents with Offline Reinforcement Learning

### v1

@@ 4.1 Experimental Setup
Offline Reinforcement Learning Dataset To train our model, we curate an offline reinforcement learning dataset derived from the Waymo Open Motion dataset (Ettinger et al., 2021). Each scene in the Waymo Open Motion dataset is autoregressively fed through the physics-enhanced Nocturne simulator to compute the per-timestep actions and factored rewards for each agent. The actions are defined by the acceleration and steering angle and the reward function is decomposed into three components: a goal position reward, a vehicle to vehicle collision reward, and a vehicle to road edge collision reward. We refer readers to Appendix D for more details on how we obtain the per-timestep actions and factored rewards. We confirm that the trajectory rollouts obtained by feeding Waymo scenes through the simulator attain a reasonable reconstruction of the ground-truth Waymo trajectories (see Table 1). Following Nocturne (Vinitsky et al., 2022), we omit bicyclist and pedestrian trajectories from the Waymo Open Motion Dataset and we omit scenes containing traffic lights. This yields a training, validation, and test set containing 134150, 9678, and 2492 scenes.
@@
Evaluation We evaluate CtRL-Sim on its ability to replicate the driving behaviours found in the Waymo Open Motion Dataset (imitation) and generate counterfactual scenes that are consistent with specified tilting coefficients (controllability). For both modes of evaluation, we use 1 second of history context and simulate an 8 second future trajectory rollout. For imitation evaluation, we evaluate on up to 8 moving agents per scene that we control with CtRL-Sim, where the remaining agents are set to log replay through physics. We evaluate on 1000 random test scenes in both modes of evaluation. Following recent work (Suo et al., 2023), we use three types of metrics for imitation evaluation: reconstruction metrics measure how well the behaviour simulation model can reconstruct the real scene, including Final Displacement Error (FDE), Average Displacement Error (ADE), and Goal Success Rate; distributional realism metrics measure the similarity of feature distributions between real and simulated scenes, summarized by the Meta-Jensen Shannon Distance (Meta-JSD) defined as the mean of the linear speed, angular speed, acceleration, and distance to nearest vehicle JSDs; and common sense metrics measure the model’s ability to avoid traffic violations, measuring Collision Rate and Offroad Rate. Refer to Appendix E for more detail on the imitation metrics.
@@
For controllability evaluation, we evaluate on 1 selected “interesting” interactive agent that is controlled by CtRL-Sim, defined as an agent who is moving and whose goal is within 10 metres of another moving agent. All agents except for the CtRL-Sim-controlled interesting agent is set to log replay through physics. We evaluate the model’s controllability through metrics aligned with the specified reward dimensions: we report the goal success rate for the goal reward control, collision rate for the vehicle-vehicle reward control, and offroad rate for the vehicle-road-edge reward control.
Reconstruction Distributional Realism Common Sense
@@
Actions-Only (Philion et al., 2023) 11.55 $\pm$ 1.19 4.74 $\pm$ 0.46 34.4 $\pm$ 1.1 14.3 $\pm$ 0.4 22.7 $\pm$ 0.8 30.0 $\pm$ 1.7
Imitation Learning 2.40 $\pm$ 0.06 1.46 $\pm$ 0.04 73.5 $\pm$ 0.9 12.2 $\pm$ 0.2 7.5 $\pm$ 0.5 13.3 $\pm$ 0.4
DT (Max Return) (Chen et al., 2021) 3.17 $\pm$ 0.15 1.66 $\pm$ 0.05 60.7 $\pm$ 0.2 12.2 $\pm$ 0.2 6.1 $\pm$ 0.7 11.6 $\pm$ 0.3
CtRL-Sim (No State Prediction) 2.51 $\pm$ 0.14 1.49 $\pm$ 0.06 66.7 $\pm$ 0.9 12.3 $\pm$ 0.3 7.6 $\pm$ 0.7 13.2 $\pm$ 0.2
CtRL-Sim (Base) 2.44 $\pm$ 0.06 1.47 $\pm$ 0.03 68.2 $\pm$ 0.8 12.3 $\pm$ 0.1 7.5 $\pm$ 0.1 13.1 $\pm$ 0.5
CtRL-Sim (Positive Tilting) 2.32 $\pm$ 0.02 1.42 $\pm$ 0.02 67.3 $\pm$ 0.6 12.2 $\pm$ 0.0 6.7 $\pm$ 0.4 12.4 $\pm$ 0.3
$\text{DT}^{*}$ (GT Initial Return) 1.93 $\pm$ 0.07 1.27 $\pm$ 0.02 72.9 $\pm$ 0.3 12.2 $\pm$ 0.1 6.6 $\pm$ 0.4 12.7 $\pm$ 0.3
@@
Table 1: Multi-agent simulation results over 1000 test scenes. This table shows how well different approaches recover the ground-truth scene. We report the mean and std across 3 seeds (training and evaluation). For each metric, the best method is bolded and the second-best method is underlined. CtRL-Sim demonstrates a good balance between common sense, realism, and reconstruction performance. denotes privileged models that require the ground-truth future trajectory.
Methods under Comparison For imitation evaluation, we compare CtRL-Sim against several relevant baselines: 1. Replay-Physics: The upper bound of imitation performance on the dataset, which employs an inverse bicycle model to obtain the ground-truth actions, and executes through the physics of the simulator, as described in Appendix D. 2. Actions-Only: An encoder-decoder model based on Philion et al. (2023), where the decoder trajectory sequences contain only actions. As the actions in CtRL-Sim are accelerations and steering rather than position offsets as used in Philion et al. (2023), there may be differences in the achieved performance. 3. Imitation Learning (IL): An identical architecture to the one described in Section 3.2 with the removal of state prediction and return information. 4. Decision Transformer: We employ two variants, where GT Initial Return specifies the initial ground-truth return-to-go from the offline RL trajectory rollout, with the goal of acting as an imitative policy. The second variant is the Max Return and follows the traditional Decision Transformer approach of selecting the maximum observable return in the dataset. The Decision Transformer architecture and training details are identical to that of CtRL-Sim except each return token precedes its corresponding state token in the sequence, and the returns and future states are not predicted by the decoder.
We evaluate multiple versions of the proposed CtRL-Sim behaviour simulation model for imitation evaluation, and we refer readers to Appendix G for details regarding the architecture and training specifications. CtRL-Sim (Base) is the base trained model without additional finetuning. CtRL-Sim (No State Prediction) is a variant of the base model where we remove the state prediction task. CtRL-Sim (Positive Tilting) applies $\kappa_{c}=10$ tilting to all components $c$ of the base CtRL-Sim model. CtRL-Sim (GT Initial Return) is similar to DT (GT Initial Return). For controllability evaluation, we evaluate on the CtRL-Sim base model and a finetuned CtRL-Sim model. The finetuned model is the base model that is finetuned on a dataset of simulated long-tail scenarios collected using the CtRL-Sim base model with tilting. This allows CtRL-Sim to be exposed to more long-tail scenarios during training, as the Waymo Open Motion dataset contains mainly nominal driving scenarios. We refer readers to Appendix I for details of our proposed finetuning procedure.
4.2 Results
Figure 4: Results of tilting with base CtRL-Sim model. We report the effects of exponential tilting with CtRL-Sim along the different reward dimensions. Left: tilting the vehicle-vehicle collision reward from -25 to 25 decreases the collision rate. Middle: tilting the vehicle-edge collision reward from -25 to 25 smoothly decreases the collision rate. Right: tilting the goal reward from -25 to 25 increases the goal success rate. We report the mean and std over three seeds.
Figure 5: Results of tilting with fine-tuned CtRL-Sim model. Finetuning yields a wider range of collision (left) and offroad (middle) behaviours (see the y axis values). Goal controllability (right) remains stable after finetuning CtRL-Sim. We report the mean and std over 3 seeds.
In Table 1, we present the multi-agent imitation results comparing the CtRL-Sim model and its variants with imitation baselines. The CtRL-Sim models perform well compared to the imitation baselines, with the CtRL-Sim (Positive Tilting) model achieving the best balance between the reconstruction metrics (1st in FDE and ADE), distributional realism (Tied 1st in Meta-JSD), and common sense metrics (2nd in Collision Rate and Offroad Rate). Although DT (Max Return) attains better collision and offroad rate, this comes at the cost of substantially worse reconstruction performance. We further conduct an ablation study confirming the importance of the future state prediction task, with CtRL-Sim (Base) outperforming CtRL-Sim (No State Prediction) across all metrics. The CtRL-Sim (Positive Tilting) model attains the best collision rate and offroad rate amongst the CtRL-Sim models, demonstrating the effectiveness of exponential tilting for steering the model towards good driving behaviours.
We emphasize that a distinctive feature of CtRL-Sim that distinguishes it from the imitation baselines in Table 1 is that it additionally enables intuitive control over the agent behaviours through exponential tilting of the return distribution. This contrasts with the Decision Transformer architecture, which, although capable of generating suboptimal behaviors by specifying low initial return-to-gos, lacks intuitive control due to the prerequisite knowledge about the return-to-go values and an absence of an intepretable mechanism for behavior modulation. On the other hand, exponential tilting of the predicted return distribution, as is done in CtRL-Sim, has a clear interpretation: negative exponential tilting yields behaviours that are worse than the average behaviours learned from the dataset, while positive exponential tilting yields better-than-average behaviours. This provides a far more intuitive interface to a practitioner who may aim to produce behaviours that are either less or more optimal than nominal driving behaviours.
Model Collision (%) Meta-JSD (×$\times 10^{-2}$)
DT (Min Return) 10.6 $\pm$ 1.6 13.4 $\pm$ 0.7
CtRL-Sim (-10 Tilt, Veh-Veh) 11.3 $\pm$ 1.2 13.2 $\pm$ 0.2
CtRL-Sim (-25 Tilt, All) 19.5 $\pm$ 2.7 15.2 $\pm$ 0.4
CtRL-Sim FT (-25 Tilt, All) 33.4 $\pm$ 2.4 20.9 $\pm$ 0.6
Table 2: CtRL-Sim more reliably generates collision scenarios than the Decision Transformer baseline, demonstrating the effectiveness of exponential tilting. Finetuning improves collision generation with CtRL-Sim. Best model is bolded and second best is underlined.
@@ 4.2 Results
We show the results of our controllability evaluation in Figure 4. For each reward dimension $c$, we exponentially tilt $\kappa_{c}$ between -25 and 25 and observe how this affects the corresponding metric of interest. In Figure 4, we observe a relatively monotonic change in each metric of interest as the tilting coefficient is increased from -25 to 25. In Figure 5, we repeat these experiments with the finetuned model. As the finetuned model has more exposure to long-tail scenarios during training, it demonstrates significant improvements over the base model in generating safety-critical driving behaviours, while demonstrating smooth changes in each metric of interest as the tilting coefficient is increased from -25 to 25. Specifically, the finetuned model observes a smooth decrease in collision rate from 31% to 11% via vehicle-vehicle exponential tilting, and a smooth decrease in offroad rate from 26% to 16% via vehicle-road-edge exponential tilting. Importantly, at -25 exponential tilting, the finetuned model is able to generate 1.9×1.9$\times$1.9 × as many collisions and offroad violations as the CtRL-Sim base model, demonstrating that finetuning CtRL-Sim on simulated long-tail scenarios improves its ability to generate long-tail counterfactual scenes. Figure 6 (and 7, 8 in Appendix ) shows qualitatively the effects of CtRL-Sim’s tilting. Table 2 compares CtRL-Sim with a Decision Transformer baseline in the generation of adversarial collision scenarios. To maximally encourage collisions, CtRL-Sim (-25 Tilt, All) applies $\kappa_{c}=-25$ tilting to all reward components $c$. This encourages collisions without constraining the behaviour to stay on onroad or reach its prescribed goal. CtRL-Sim FT (-25 Tilt, All) applies the same tilting to the finetuned model. We also compare with CtRL-Sim (-10 Tilt, Veh-Veh) where only the vehicle-vehicle reward component has $-10$ tilting applied. We compare with the Decision Transformer baseline (DT (Min Return)) where the target return is specified as the minimum possible return across the three reward dimensions. We follow the same evaluation protocol as in Figure 4. Table 2 shows that the CtRL-Sim models produce significantly higher collision rates than the Decision Transformer baseline, with the finetuned model producing a collision in $1/3$ of the scenes. Moreover, the CtRL-Sim (-10 Tilt, Veh-Veh) model attains a higher collision rate than DT (Min Return) while achieving a lower Meta-JSD, highlighting the effectiveness of CtRL-Sim in generating realistic collision behaviours.
@@
Figure 6: Qualitative results of vehicle-vehicle and vehicle-edge tilting. We show the evolution of four traffic scenes with the top panels applying positive tilting to the CtRL-Sim-controlled agent (shown in teal) and the bottom panels applying negative tilting to the same CtRL-Sim-controlled agent. Bounding boxes outlined in red contain a traffic violation. All other agents are set to log-replay through physics, with the agent interacting with the CtRL-Sim-controlled agent denoted in pink. Goals are denoted by small circles. Exponential tilting enables control over the generation of realistic vehicle-vehicle and vehicle-edge collision scenarios.

### v2

@@ 3.1 Experimental Setup
Offline Reinforcement Learning Dataset To train our model, we curate an offline reinforcement learning dataset derived from the Waymo Open Motion dataset [14]. We first extend Nocturne by integrating a physics engine based on the Box2D library for enabling realistic vehicle dynamics and collisions, detailed in Appendix C.1. Each scene in the Waymo dataset is fed through the physics-enhanced Nocturne simulator to compute the per-timestep actions and factored rewards for each agent. Refer to Appendix C for more details regarding the offline RL dataset collection.
@@
Evaluation We evaluate CtRL-Sim on its ability to replicate the driving behaviours found in the Waymo Open Motion Dataset (imitation) and generate counterfactual scenes that are consistent with specified tilting coefficients (controllability). For both modes of evaluation, we use 1 second of history and simulate an 8 second future rollout. For imitation, we evaluate on up to 8 moving agents per scene that we control with CtRL-Sim, where the remaining agents are set to log replay through physics. We evaluate on 1000 random test scenes in both modes of evaluation. Following recent work [7], we use three types of metrics for imitation evaluation: reconstruction metrics, such as Final Displacement Error (FDE), Average Displacement Error (ADE), and Goal Success Rate; a distributional realism metric (JSD) defined by the mean of the Jensen-Shannon Distances computed on linear speed, angular speed, acceleration, and distance to nearest vehicle features between real and simulated scenes; and common sense metrics measured by Collision and Offroad rate.
@@
For controllability evaluation, we evaluate on 1 selected “interesting” interactive agent that is controlled by CtRL-Sim, defined as an agent who is moving and whose goal is within 10 metres of another moving agent. All agents except for the CtRL-Sim-controlled interesting agent are set to log replay through physics. We evaluate the model’s controllability through metrics aligned with the specified reward dimensions: we report the goal success rate for the goal reward control, collision rate for the vehicle-vehicle reward control, and offroad rate for the vehicle-road-edge reward control.
ADE FDE Goal Success JSD Collision Off Road Per Scene
Method (m) $\downarrow$ (m) $\downarrow$ Rate (%) $\uparrow$ (×$\times 10^{-2}$) $\downarrow$ (%) $\downarrow$ (%) $\downarrow$ Gen. Time (s) $\downarrow$
Replay-Physics $0.47$ $0.97$ $87.3$ $7.6$ $2.8$ $10.7$ $1.1$
Actions-Only [9] $4.81$$\pm$0.52 $11.89$$\pm$1.42 $32.7$$\pm$1.4 $10.4$$\pm$0.3 $19.9$$\pm$1.2 $27.6$$\pm$1.0 $3.3$
Imitation Learning $1.24$$\pm$0.05 $1.95$$\pm$0.10 $77.4$$\pm$1.3 $8.3$$\pm$0.1 $5.8$$\pm$0.2 $12.1$$\pm$0.2 $3.4$
DT (Max Return) [15] $1.56$$\pm$0.04 $3.07$$\pm$0.16 $63.3$$\pm$0.8 $8.4$$\pm$0.1 $5.3$$\pm$0.3 $11.0$$\pm$0.2 $20.7$
CTG++ [11] $1.73$$\pm$0.10 $4.02$$\pm$0.32 $38.8$$\pm$5.4 $7.4$$\pm$0.2 $5.9$$\pm$0.4 $15.0$$\pm$1.5 $44.0$
CtRL-Sim (No State Prediction) $1.32$$\pm$0.03 $2.21$$\pm$0.06 $72.4$$\pm$0.8 $8.2$$\pm$0.2 $6.1$$\pm$0.4 $12.0$$\pm$0.3
CtRL-Sim (Base) $1.29$$\pm$0.04 $2.13$$\pm$0.08 $73.0$$\pm$1.3 $8.1$$\pm$0.2 $5.8$$\pm$0.4 $11.8$$\pm$0.2 $8.2$
CtRL-Sim (Positive Tilting) $1.25$$\pm$0.03 $2.04$$\pm$0.08 $72.9$$\pm$1.5 $7.9$$\pm$0.1 $5.3$$\pm$0.2 $11.0$$\pm$0.2
DT (GT Initial Return) $1.10$$\pm$0.02 $1.58$$\pm$0.07 $77.5$$\pm$1.5 $8.4$$\pm$0.1 $5.3$$\pm$0.3 $11.9$$\pm$0.3 $20.8$
CtRL-Sim (GT Initial Return) $1.09$$\pm$0.02 $1.60$$\pm$0.06 $77.2$$\pm$1.1 $8.1$$\pm$0.2 $5.6$$\pm$0.4 $12.2$$\pm$0.1
@@
Methods under Comparison For imitation evaluation, we compare CtRL-Sim against several relevant baselines: 1. Replay-Physics employs an inverse bicycle model to obtain the ground-truth log-replay actions and executes through the simulator. 2. Actions-Only is an encoder-decoder model inspired by [9] where the decoder trajectory sequences only contain actions. 3. Imitation Learning (IL) is identical to the architecture in Section 2.2 except with the removal of returns and the future state prediction. 4. Decision Transformer (DT): The GT Initial Return variant specifies the initial ground-truth return-to-go from the offline RL dataset, with the goal of acting as an imitative policy. Max Return follows the standard DT approach of selecting the maximum observable return in the dataset. The DT architecture is identical to that of CtRL-Sim except the return token precedes the state token, and the returns and future states are not predicted by the decoder. 4. CTG++ is a reimplementation of [11], a competitive Transformer-based diffusion model for behaviour simulation.
We evaluate the following variants of the proposed CtRL-Sim model: 1. CtRL-Sim (Base) is the CtRL-Sim model trained on the offline RL dataset. 2. CtRL-Sim (No State Prediction) is the base model trained without the state prediction task. 3. CtRL-Sim (Positive Tilting) applies $\kappa_{c}=10$ tilting to all components $c$ of the base model. 4. CtRL-Sim (GT Initial Return) is similar to DT (GT Initial Return). For controllability evaluation, we evaluate on the CtRL-Sim base model and a finetuned CtRL-Sim model (CtRL-Sim FT). The finetuned model takes a trained base model and finetunes it on a dataset of simulated long-tail scenarios that we collect using an existing simulated collision generation method CAT [21]. This allows CtRL-Sim to be exposed to more long-tail collision scenarios during training, as the Waymo Open Motion dataset mainly contains nominal driving. We refer readers to Appendix H for details of CAT and our proposed finetuning procedure.
3.2 Results
In Table 1, we present the multi-agent imitation results comparing the CtRL-Sim model and its variants with imitation baselines. The CtRL-Sim models perform competitively with the imitation baselines, with the CtRL-Sim (Positive Tilting) model achieving a good balance between distributional realism (2nd in JSD), reconstruction performance (2nd in FDE, ADE), common sense (Tied 1st in Collision and Offroad Rate), and efficiency (5.4×5.4$\times$5.4 × faster than CTG++). Although DT (Max Return) attains equal collision and offroad rates as CtRL-Sim, this comes at the cost of substantially worse reconstruction performance. We further validate the importance of the future state prediction task, with CtRL-Sim (Base) outperforming CtRL-Sim (No State Prediction) across all metrics. The CtRL-Sim (Positive Tilting) model attains the best collision rate and offroad rate, demonstrating the effectiveness of exponential tilting for steering the model towards good driving behaviours.
@@ 3.2 Results
We emphasize that a distinctive feature of CtRL-Sim that distinguishes it from the imitation baselines in Table 1 is that it additionally enables intuitive control over the agent behaviours through exponential tilting of the return distribution. This contrasts with DT, which, although capable of generating suboptimal behaviours by specifying low initial return-to-gos, lacks intuitive control due to the prerequisite knowledge about the return-to-go values and an absence of an interpretable mechanism for behaviour modulation. By constrast, the exponential tilting of the predicted return distribution employed in CtRL-Sim has a clear interpretation: negative exponential tilting yields behaviours that are worse than the average behaviours learned from the dataset, while positive exponential tilting yields better-than-average behaviours. This provides a more intuitive interface to a practitioner who may aim to produce behaviours that are either less or more optimal than nominal driving behaviours.
Planner Metrics Adversary Realism
Progress Coll. w/ Adv. JSD Coll. Speed
Adv. Method Tilt Reactive? Control? (m) $\downarrow$ (%) $\uparrow$ (×$\times 10^{-2})$downarrow (m/s) $\downarrow$
CAT $53.3$ $61.4$ $18.7$ $6.9$
CtRL-Sim $-10$ $57.5$$\pm$0.1 $10.0$$\pm$0.5 $10.8$$\pm$0.3 $7.4$$\pm$0.5
$10$ $57.7$$\pm$0.1 $8.7$$\pm$0.5 $10.3$$\pm$0.5 $8.3$$\pm$0.6
$-10$ $56.1$$\pm$0.2 $33.8$$\pm$1.9 $17.4$$\pm$0.6 $6.3$$\pm$0.2
CtRL-Sim FT $10$ $57.1$$\pm$0.1 $18.5$$\pm$1.6 $12.6$$\pm$0.7 $6.1$$\pm$0.2
$50$ $57.4$$\pm$0.2 $12.8$$\pm$0.2 $15.6$$\pm$1.2 $6.0$$\pm$0.3
@@
We show the results of our controllability evaluation in Figure 4. For each reward dimension $c$, we exponentially tilt $\kappa_{c}$ between -25 and 25 and observe how this affects the corresponding metric of interest. We also show the results of DT when conditioning on the minimum and maximum possible return. For both the base and finetuned CtRL-Sim models, we observe a relatively monotonic change in each metric of interest as the tilting coefficient is increased from -25 to 25. As the finetuned model is exposed to collision scenarios during finetuning, it demonstrates significant improvements over the base model in generating bad driving behaviours. Specifically, at -25 tilting, the finetuned model is able to generate 2.1×2.1$\times$2.1 × as many collisions and 1.8×1.8$\times$1.8 × as many offroad violations as the base model. Figure 5 (and 7, 8 in Appendix J) shows qualitatively the effects of tilting.
Table 2 evaluates CtRL-Sim’s ability to produce adversarial agents that collide with a data-driven planner. We evaluate on a held-out test set of two-agent interactive scenarios from the Waymo dataset, where one interacting agent is controlled by the planner and the other is controlled by the adversary. We use a positively-tilted CtRL-Sim base model as our planner, due to its demonstrated ability to produce good driving behaviours in Table 1. For adversarial scenario generation, we compare the base CtRL-Sim model against the CtRL-Sim FT model. With -10 tilting applied, the finetuned model generates 238 more collisions with the planner than the base model over 1000 scenes, which we attribute to its exposure to simulated collision scenarios during finetuning. Notably, CtRL-Sim FT was finetuned in only 30 minutes on 1 NVIDIA A100-Large GPU and only 3500 CAT scenarios. This underscores CtRL-Sim’s capability to flexibly incorporate data from various sources through finetuning, thereby enabling the generation of new kinds of driving behaviours. Importantly, after finetuning, CtRL-Sim FT largely retains its ability to produce good driving behaviours. This is evidenced by a 21.1 percentage point decrease in the planner’s collision rate when using a +50 positively tilted finetuned model as the adversary.
In Table 2, we also compare CtRL-Sim against a state-of-the-art collision generation method, CAT [21], which uses a goal-conditioned trajectory forecasting model to select plausible adversarial trajectories that overlap with the ego plan. Although CAT generates more collisions with the planner, CAT is not controllable in that it cannot control the degree to which the agents are adversarial, which is a distinguishing feature of CtRL-Sim. Furthermore, CAT agents are non-reactive to the ego’s actions as the trajectory is fixed at the beginning of the simulation, which severely limits the realism of agents controlled by CAT. This is evidenced by a larger adversary collision speed than all finetuned CtRL-Sim models and is also validated qualitatively in the attached supplementary videos. As collision realism is hard to quantitatively assess, we further conduct a user study to confirm that CtRL-Sim adversarial scenarios are indeed more realistic than CAT adversarial scenarios, with details and results reported in Appendix I.
@@
Figure 5: Qualitative results of vehicle-vehicle and vehicle-edge tilting. Two traffic scenes comparing positive tilting of the CtRL-Sim-controlled agent (shown in teal) with negative tilting for the same agent. Bounding boxes in red indicate traffic violations. Other agents log-replay through physics, with interacting agents in pink. Goals are marked by small circles.